Suppose that you are trying to decide whether to spend $1,000 on stock issued by WildWeb or on a bond issued by the same company. There is a 40 percent chance that the value of the stock will rise to $2,200 at the end of the year and a 60 percent chance that the stock will be worthless at the end of the year. The bond promises an interest rate of 30 percent per year, and it is certain that the bond and interest ill be repaid at the end of the year Assuming that your time horizon is exactly one year, will you choose the słock or the bond? By how much is your expected end-of-year wealth reduced if you make the wrong choioe? $ Suppose the odds of success improve for WidWeb: Now thore is a 50 percent chance that the value of the stook will be $2,200 at years end and only a 50 percent chance that it will be worthless. Should you now choose the stock or the bond?
Answer #1

Expected end of year wealth of investing in stock= 2200* 0.4 + 0* 0.6 = $ 880

Expected end of year wealth of investing in bond= (1000*1.3) * 1 = $ 1300

As the expected end of year wealth is higher by investing in bonds when compared to the expected end of year wealth by investing in stocks, we will choose to invest in bonds.

If a wrong choice is made, expected end of year wealth will decrease by ($1300 - $880) = $ 420

Now, if the probability of $1000 becoming $2200 by investing in stocks becomes 50%

Expected end of year wealth of investing in stock= 2200* 0.5 + 0* 0.5 = $ 1100

Expected end of year wealth of investing in bond= (1000*1.3) * 1 = $ 1300

As the expected end of year wealth is higher by investing in bonds when compared to the expected end of year wealth by investing in stocks, we will again choose to invest in bonds.